# Implementation and characterization of the dice lattice in the electron quantum simulator

### v2

@@ I Introduction
In the past few years, there has been a growing interest in studying lattice systems characterized by spectra containing one or more quasi-flat bands [1]. This interest stems from the discovery of superconductivity and other correlated phases in layered systems [2, 3]. The prevailing understanding suggests that correlated phenomena are observable in lattice systems with quasi-flat bands, primarily due to the careful comparison of energy scales. Specifically, as electronic kinetic energy diminishes to the same order of magnitude as, or even below, the interaction energy between electrons, correlation effects become visible. This observation stems from a fundamental competition between kinetic and interaction energies within these systems, highlighting the intricate interplay of energy scales at play.
@@
Here, we will investigate and propose a possible experimental implementation of another lattice with a flat band at zero energy: the dice lattice [5]. We present a sketch of the dice lattice in Figure 1(a). It is a triangular lattice with a base of the sites in the unit cell; as in the case of the Lieb lattice, the connectivity of the three sites is not homogeneous. Specifically, we have one site, dubbed H with connectivity 6, and two inequivalent sites, named R and R’ with connectivity 3. Similar to the Lieb lattice, the energy spectrum of the dice lattice consists of three bands, two of which are energy symmetric, and the third one is completely flat and at zero energy. The three bands have an accidental degeneracy at the K point of the BZ. In this point, it is possible to describe the long-wavelength low-energy expansion of the system Hamiltonian in terms of an integer pseudo-spin theory [29]. Some previous work proposed the experimental implemenation the dice lattice in the framework of cold atoms in optical lattices [29, 30].
@@
In this article, we propose a realization of the dice lattice in the electron quantum simulator. Specifically, we will consider the case of decorating with CO molecules the surface (111) of Cu [18, 15]. We propose in total three different realizations differing for the amount of CO molecules implemented and the area of the unit cell. We evaluate the spectral properties of these configurations and show how they map into the simplest next-nearest-neighboring tight-binding approximation. The advantage of implementing the dice lattice compared to the Lieb lattice roots is down to the compatibility between the space group of the dice lattice with one of the Cu(111) substrate.
The paper is organized as follows. In Section II.1, we revise the properties of the dice lattice spectrum in the next-nearest-neighboring tight-binding approximation. In Section II.2, we present our three proposals for the experimental implementation and find the corresponding parameters to map into next-nearest-neighboring tight-binding models. In Section II.3, we revise the long-wavelength low-energy approximation for the dice lattice Hamiltonian expanded around the K point of the BZ, including terms describing the next-nearest-neighboring hopping terms. Finally, in Sections II.4, we revise the properties of the dice lattice in a perpendicular magnetic field in the tight-binding approximation. In Sections III, we summarize the work and give some possible outlook.
@@ II Results
In the following, we describe the general properties of the dice lattice within the simplest tight binding model without a magnetic field. Successively, we present a possible theoretical implementation with an electron quantum simulator platform based on CO molecules on Cu(111). The numerical implementation is based on the nearly-free electron model, in which a muffin-tin potential models the CO molecules, and we compare the results with a next-nearest-neighbor tight-binding model. Finally, we present the results for the spectrum when a perpendicular magnetic field is applied.
@@ II.1 Tight-binding model
The dice lattice is a regular tesselation of the Euclidean plane with regular $\pi/3$ rhombi, resulting in a bipartite structure with hexagonal symmetry. Each unit cell contains three sites; in the following, we name them a hub, indicated with the letter H, and two rims, indicated with R and R’, respectively [31]. We present a sketch of the lattice in Figure 1(a). The following translational vectors characterize the unit cell:
@@
where $a$ is the distance between the lattice sites. In the following, we consider a rhombic unit cell for simplicity, while the Wigner-Seitz cell is hexagonal. We describe the dice lattice with a simple $s$-orbital tight-binding model, including next-neighbor and next-nearest-neighboring hopping terms, $t$ and $t^{\prime}$, respectively. Neglecting the next-nearest-neighboring hopping terms $t^{\prime}$, the hub sites are sixfold connected, while the rim sites are threefold connected.
@@
$\hat{h}(\bm{k})=\begin{pmatrix}0&&f_{\text{RH}}(\bm{k})\,t&&f_{\text{RR'}}(\bm{k})\,t^{\prime}\\ f_{\text{RH}}^{\ast}(\bm{k})\,t&&-\Delta&&f_{\text{HR'}}(\bm{k})\,t\\ f_{\text{RR'}}^{\ast}(\bm{k})\,t^{\prime}&&f_{\text{HR'}}^{\ast}(\bm{k})\,t&&0\end{pmatrix}.$ (5)
This Hamiltonian is expressed in the following base $\Psi=(\psi_{\text{R}},\psi_{\text{H}},\psi_{\text{R'}})^{\text{T}}$ and we have introduced the following auxiliaries functions:
@@
here, $\bm{k}$ is the crystal momentum vector defined within the BZ. In Eq. (5), we have introduced an energy term $\Delta$ describing the difference between the hub and rim on-site energy. It is not possible to find a compact expression for the energy bands; however, for $t^{\prime}=0$, the energy spectrum is given by
@@
We start presenting the spectral properties of the dice lattice in the simple case of $t^{\prime}=\Delta=0$. In this limit, of the three bands composing the spectrum, two are dispersive $\mathcal{E}_{\pm}$, and one is completely flat at zero energy $\mathcal{E}_{\text{c}}$igures 1(c). The two dispersive bands are identical to the one of the honeycomb lattice [38]. The flat band $\mathcal{E}_{\text{c}}$ arises because of the bipartite nature of the dice lattice [32, 8] and can be interpreted in terms of CLSs as well [39]. The three bands are degenerate in the K and K’ points of the BZ. This threefold degeneracy is accidental since, in two dimensions, there are no irreducible representations of order three [40, 41]. Performing a low-energy long-wavelength approximation of the energy spectrum around this K or K’, it is possible to obtain an effective pseudo-spin one Hamiltonian model for the system [29, 30, 28, 42], see Section II.3 for additional details.
By inspecting Eq. (7b), we notice that the onsite energy $\Delta$ shifts the two dispersive bands and partially removes the degeneracy at the K(K’)-points by opening a gap between a dispersive band and the flat one — Figure 2(a). Changing the sign of $\Delta$, the energy gap opens between the flat band and the other dispersive one. Furthermore, a finite value of the next-nearest-neighboring hopping $t^{\prime}$ induces a finite dispersion in the flat band — Figure 2(b), but it does not open an energy gap at the three-fold degeneracy point in $K$. The energy symmetry is broken when $\Delta$ and $t^{\prime}$ are finite — see Figure 1(d). We note on passing that the effects of $t^{\prime}$ are similar to the effects of evaluating the energy spectrum by considering an overlap integral [43].
In Figures 2(b) and 2(d), we present the density of states (DOS). corresponding to the spectra depicted in Figures 2(a) and 2(c), respectively. We define the DOS as
@@ II.2 Implementation with CO molecules on Cu(111)
where $m^{*}$ is the effective mass of the free electrons on Cu(111) [22], the periodic potential $V_{\text{CO}}(\bm{R})$ is obtained by considering the set of all CO molecules. Without loss of generality, we can assume that each CO molecule at the position $\bm{r}_{0}$ is a cylindrical repulsive potential for the Cu(111) surface electrons:
$V(\bm{r}-\bm{r}_{0})=V_{0}\begin{cases}1&|\bm{r}-\bm{r}_{0}|\leqslant D\\ 0&|\bm{r}-\bm{r}_{0}|>D\end{cases}.$ (10)
The parameters of this potential are chosen in agreement with the experiments as $V_{0}=0.9\,\text{eV}$ and a cylinder diameter of $D=0.6\,\text{nm}$ [18, 25, 22]. The complete set of CO molecules can be described by a periodic potential $V_{\text{CO}}(\bm{R})$ that can be expanded in terms of Fourier components as
@@
where $\mathcal{J}_{1}(x)$ are the Bessel functions of the first order. As a consequence summing up over all the Fourier components $V_{\bm{G}}$, we can rewrite Eq. (11) as:
@@
$=\frac{2\pi r_{0}V_{0}}{A}\sum_{\{\ell,m\}\in\mathbb{Z}}\frac{\mathcal{J}_{1}(|\ell\bm{b}_{1}+m\bm{b}_{2}|D)}{|\ell\bm{b}_{1}+m\bm{b}_{2}|}\text{e}^{\text{i}(\ell\bm{b}_{1}+m\bm{b}_{2})\cdot\bm{R}}$
@@
At this point, assuming that the system wave function is also periodic [44], we can transform the Schrördinger equation in a set of coupled linear equations
@@
[…]
@@
In Figures 4(d), 4(e), and 4(f), we present the LDOS $\rho(\bm{r},E)$ at the center of the rim and hub sites for the three configurations. These results help us to understand the role played by the next-nearest-neighboring model ($t^{\prime}\neq 0$). As introduced before, the perfectly flat band case $t^{\prime}=0$ can be characterized by a linear combination of CLSs around the hub sites [5, 8, 46, 39]. Surprisingly, for $t^{\prime}\neq 0$, we can observe that the LDOS is still localized around the rim sites for energies corresponding to the $\mathcal{E}_{\text{c}}$ band — Figure 2(b) and  2(d). The survival of imperfect CLSs is more evident from the plots of the LDOS in space Figures 4(g), 4(h), and 4(i), these density plots for $\rho(\bm{r},E)$ are obtained for an energy corresponding crossing between the upper and the central band at the K point. Here, it is possible to recognize that the LDOS still has a shape resembling the CLSs around the hub sites, with values that are much larger on the rim sites compared to the hub sites [8].
@@ II.3 Low-energy theory
In this section, we will present a low-energy, long-wavelength approximation for the Hamiltonian (5) around the K/K’ points. In the limit of $t^{\prime}=0$, the approximated Hamiltonian reads
@@
$\Sigma_{x}$ $=\frac{1}{\sqrt{2}}(\lambda_{1}+\lambda_{6})=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ (18d)
$\Sigma_{y}$ $=\frac{1}{\sqrt{2}}(\lambda_{2}+\lambda_{7})=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-\text{i}&0\\ \text{i}&0&-\text{i}\\ 0&\text{i}&0\end{pmatrix}\!\!,$ (18h)
$\ \Sigma_{z}$ $=(\lambda_{3}+\sqrt{3}\lambda_{8})=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix},$ (18l)
@@
that is unitary $\Gamma^{\dagger}\Gamma=\Gamma^{2}=\mathbb{I}_{3}$ where $\mathbb{I}_{3}$ is the identity matrix of dimension 3, and anticommutes with the Hamiltonian (17): $\{\Gamma,\mathcal{H}_{0}\}=0$. This chiral symmetry operator governs the energy symmetry of the system, see Fig. 1(c)  [^1] Here the chiral symmetry is simply anticommuting with the system Hamiltonian since the flat band is at zero energy. In other lattice systems with three sites in the unit cell, as the kagome lattice, the flat band is at finite energy. Thus they require the introduction of a generalized chiral symmetry [56, 57, 58]..
@@
We can complement the matrices (II.3) with two additional Gell-Mann matrices, $\lambda_{4}$ and $\lambda_{5}$, so to describe the corresponding low-energy next-nearest-neighboring hopping term $t^{\prime}$. This Hamiltonian term reads:
@@ II.4 Magnetic field properties: tight-binding approach
$\mathcal{M}$ $=\frac{t}{4}\begin{pmatrix}0&\text{e}^{-\frac{1}{2}\,\text{i}\,\pi f(6\,n_{1}-1)}&0\\ \text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}-1)}&-\frac{\Delta}{t}&\text{e}^{-\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}\\ 0&\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}&0\end{pmatrix},$ (27d)
$\mathcal{P}_{1}$ $=\begin{pmatrix}0&0&0\\ t&0&0\\ t^{\prime}\,\text{e}^{\frac{3}{2}\,\text{i}\,\pi\,f\,(2\,n_{1}+1)}&t&0\end{pmatrix},$ (27h)
$\mathcal{P}_{2}$ $=\begin{pmatrix}0&t\,\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}-1)}&t^{\prime}\\ 0&0&t\,\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}\\ 0&0&0\end{pmatrix},$ (27l)
$\mathcal{P}_{3}$ $=\begin{pmatrix}0&0&0\\ 0&0&0\\ t^{\prime}\,\text{e}^{-\frac{3}{2}\,\text{i}\,\pi\,f\,(2\,n_{1}+1)}&0&0\end{pmatrix}$ (27p)
contain the magnetic hopping terms. We have omitted, for simplicity in Eq. (27d), a constant shift $\epsilon$ corresponding to the effective shift of the tight-binding model fitted over the three experimental proposals.
Typically, for the dice lattice, the magnetic flux is measured through an elementary rhombus (one-third of the unit cell surface) and normalized with respect to the $\phi_{0}=h/e$ magnetic flux quantum, that is $f=B\,\sqrt{3}\,a^{2}/(2\,\phi_{0})$. The calculation of the spectrum as a function of $f$ gives rise to the Hofstadter’s butterflies [51] presented in Figsures 6 and 7. In particular, Figure 6(a) represents the ideal nearest neighbor $t=1$, $t^{\prime}=0$ case [31], whereas in Figure 6(b), we present the case of the triangular lattice corresponding to the $t=1$, $t^{\prime}=-t$ case [52]. Interestingly, for the case of the dice lattice at $f=1/2$, all the states collapse in only three energy eigenvalues. This is a consequence of the Aharonov-Bohm caging mechanism [31], for which the number of sites visited by an initially localized wave packet is limited due to destructive interferences, resulting in an extreme localization phenomenon.

### v3

@@ I Introduction
In the past few years, there has been a growing interest in studying lattice systems characterized by spectra containing one or more quasi-flat bands [1]. This interest stems from the discovery of superconductivity and other correlated phases in layered systems [2, 3]. The prevailing understanding suggests that correlated phenomena are observable in lattice systems with quasi-flat bands, primarily due to the careful comparison of energy scales. Specifically, correlation effects become visible as electronic kinetic energy diminishes to the same order of magnitude as, or even below, the electron interaction energy, thus highlighting an intricate interplay of energy scales at play.
@@
Here, we will investigate and propose a possible experimental implementation of another lattice with a flat band at zero energy: the dice lattice [5]. We present a sketch of the dice lattice in Figure 1(a). It is a triangular lattice with a base of three sites in the unit cell; as in the case of the Lieb lattice, the connectivity of the three sites is not homogeneous. Specifically, we have one site, dubbed H with connectivity 6, and two inequivalent sites, named R and R’ with connectivity 3. Similar to the Lieb lattice, the energy spectrum of the dice lattice consists of three bands, two of which are energy symmetric, and the third one is completely flat and at zero energy. The three bands have an accidental degeneracy at the K point of the BZ. In this point, it is possible to describe the long-wavelength low-energy expansion of the system Hamiltonian in terms of an integer pseudo-spin theory [29]. Some previous work proposed the experimental implementation the dice lattice in the framework of cold atoms in optical lattices [29, 30].
@@
In this article, we propose a realization of the dice lattice in the electron quantum simulator. Specifically, we will consider the case of decorating with CO molecules the surface (111) of Cu [18, 15]. We propose in total three different realizations differing for the amount of CO molecules implemented and the area of the unit cell. We evaluate the spectral properties of these configurations and show how they map into the simplest next-nearest-neighboring tight-binding approximation. The advantage of implementing the dice lattice compared to the Lieb lattice roots is down to the compatibility between the space group of the dice lattice with one of the Cu(111) substrates. It is important to note that the set of parameters we used for our implementation are in agreement with the experimental parameters [18, 25, 38, 20, 22].
The paper is organized as follows. In Section II.1, we revise the properties of the dice lattice spectrum in the next-nearest-neighboring tight-binding approximation. In Section II.2, we present our three proposals for the experimental implementation and find the corresponding parameters to map into next-nearest-neighboring tight-binding models. In Section II.3, we revise the long-wavelength low-energy approximation for the dice lattice Hamiltonian expanded around the K point of the BZ, including terms describing the next-nearest-neighboring hopping terms. Finally, in Sections II.4, we revise the properties of the dice lattice in a perpendicular magnetic field in the tight-binding approximation. In Sections III, we summarize the work and give some possible outlook. In Appendix A, we show what is happening when changing the system parameters compared to the standard one used in the experiments [18, 25, 38, 20, 22].
@@ II Results
In the following, we describe the general properties of the dice lattice within the simplest tight-binding model without a magnetic field. Successively, we present a possible theoretical implementation with an electron quantum simulator platform based on CO molecules on Cu(111). The numerical implementation is based on the nearly-free electron model, in which a muffin-tin potential models the CO molecules. We compare the results with a next-nearest-neighbor tight-binding model. Finally, we present the results for the spectrum when a perpendicular magnetic field is applied.
@@ II.1 Tight-binding model
The dice lattice is a regular tessellation of the Euclidean plane with regular $\pi/3$ rhombi, resulting in a bipartite structure with hexagonal symmetry. Each unit cell contains three sites; in the following, we name them a hub, indicated with the letter H, and two rims, indicated with R and R’, respectively [31]. We present a sketch of the lattice in Figure 1(a). The following translational vectors characterize the unit cell:
@@
where $a$ is the distance between the lattice sites. In the following, we consider a rhombic unit cell for simplicity, while the Wigner-Seitz cell is hexagonal. We describe the dice lattice with a simple $s$-orbital tight-binding model, including next-neighbor (solid lines in Figure 1(a)) and next-nearest-neighboring (dashed lines in Figure 1(a)) hopping terms, $t$ and $t^{\prime}$, respectively. Neglecting the next-nearest-neighboring hopping terms $t^{\prime}$, the hub sites are sixfold connected, while the rim sites are threefold connected. Thus, including $t^{\prime}$ terms will destroy this property, and all the lattice sites will have the same connectivity. This is important since the flat band originates in unequal connectivity among the sites.
@@
$\hat{h}(\bm{k})=\begin{pmatrix}0&&f_{\text{RH}}(\bm{k})\,t&&f_{\text{RR'}}(\bm{k})\,t^{\prime}\\ f_{\text{RH}}^{\ast}(\bm{k})\,t&&-\Delta&&f_{\text{HR'}}(\bm{k})\,t\\ f_{\text{RR'}}^{\ast}(\bm{k})\,t^{\prime}&&f_{\text{HR'}}^{\ast}(\bm{k})\,t&&0\end{pmatrix}$ (2)
expressed in the following base $\Psi=(\psi_{\text{R}},\psi_{\text{H}},\psi_{\text{R'}})^{\text{T}}$. Here, we have introduced the following auxiliaries functions:
@@
where $\bm{k}$ is the crystal momentum vector defined within the BZ, see Figure 1(b). In Eq. (2), we have introduced an energy term $\Delta$ describing the difference between the hub and rim on-site energy. It is not possible to find a compact expression for the energy bands; however, for $t^{\prime}=0$, the energy spectrum is given by
@@
We start presenting the spectral properties of the dice lattice in the simple case of $t^{\prime}=\Delta=0$. In this limit, of the three bands composing the spectrum, two are dispersive $\mathcal{E}_{\pm}$, and one is completely flat at zero energy $\mathcal{E}_{\text{c}}$igures 1(c). The two dispersive bands are identical to the one of the honeycomb lattice [39]. The flat band $\mathcal{E}_{\text{c}}$ arises because of the bipartite nature of the dice lattice [32, 8] and can be interpreted in terms of CLSs as well [40]. The three bands are degenerate in the K and K points of the BZ. This threefold degeneracy is accidental since, in two dimensions, there are no irreducible representations of order three [41, 42]. Performing a low-energy long-wavelength approximation of the energy spectrum around this K or K, it is possible to obtain an effective pseudo-spin one Hamiltonian model for the system [29, 30, 28, 43], see Section II.3 for additional details.
By inspecting Eq. (4b), we notice that the onsite energy $\Delta$ shifts the two dispersive bands and partially removes the degeneracy at the K(K)-points by opening a gap between a dispersive band and the flat one — Figure 2(a). Changing the sign of $\Delta$, the energy gap opens between the flat band and the other dispersive one. Thus, we conclude that this term breaks the three-fold degeneracy at the K(K)-points. Furthermore, a finite value of the next-nearest-neighboring hopping $t^{\prime}$ induces a finite dispersion in the flat band — Figure 2(b), but it does not open an energy gap at the three-fold degeneracy point in $K$. The energy symmetry is broken when $\Delta$ and $t^{\prime}$ are finite — see Figure 1(d). We note on passing that the effects of $t^{\prime}$ are similar to the effects of evaluating the energy spectrum by considering an overlap integral [44].
In Figures 2(b) and 2(d), we present the density of states (DOS) corresponding to the spectra depicted in Figures 2(a) and 2(c), respectively. We define the DOS as
@@ II.2 Implementation with CO molecules on Cu(111)
where $m^{*}$ is the effective mass of the free electrons on Cu(111) [22], the periodic potential $V_{\text{CO}}(\bm{R})$ is obtained by considering the set of all CO molecules. Without loss of generality, we can assume that each CO molecule at the position $\bm{r}_{0}$ is a cylindrical repulsive potential for the Cu(111) surface electrons [18]:
$V(\bm{r}-\bm{r}_{0})=V_{0}\begin{cases}1&|\bm{r}-\bm{r}_{0}|\leqslant\frac{D}{2}\\ 0&|\bm{r}-\bm{r}_{0}|>\frac{D}{2}\end{cases}.$ (7)
The parameters of this potential are chosen in agreement with the experiments as $V_{0}=0.9\,\text{eV}$ and a cylinder diameter of $D=0.6\,\text{nm}$ [18, 25, 22]. In Appendix A, we show what happens by changing the shape of the potential describing the CO molecules and the behavior by changing the system parameters away from the optimal ones. The complete set of CO molecules can be described by a periodic potential $V_{\text{CO}}(\bm{R})$ that can be expanded in terms of Fourier components as
@@
where $\mathcal{J}_{1}(x)$ are the Bessel functions of the first order, and the integral is performed over the unit cell (UC). Summing up over all the Fourier components $V_{\bm{G}}$, we can rewrite Eq. (8) as:
@@
$=\frac{2\pi r_{0}V_{0}}{A}\!\!\!\!\sum_{\{\ell,m\}\in\mathbb{Z}}\!\!\frac{\mathcal{J}_{1}(|\ell\bm{b}_{1}+m\bm{b}_{2}|D)}{|\ell\bm{b}_{1}+m\bm{b}_{2}|}\text{e}^{\text{i}(\ell\bm{b}_{1}+m\bm{b}_{2})\cdot\bm{R}}.$
@@
At this point, assuming that the system wave function is also periodic [45], we can transform the Schrödinger equation in a set of coupled linear equations
@@
$\epsilon/t$ $\Delta/t$ $t^{\prime}/t$${}_{\text{CO}}$ unit cell $\square$ (nm2)
@@
In Figures 4(d), 4(e), and 4(f), we present the LDOS $\rho(\bm{r},E)$ at the center of the rim and hub sites for the three configurations, this quantity would be accessible via STM measurements. These results help us to understand the role played by the next-nearest-neighboring model ($t^{\prime}\neq 0$). As introduced before, the perfectly flat band case $t^{\prime}=0$ can be characterized by a linear combination of CLSs around the hub sites [5, 8, 47, 40]. Surprisingly, for $t^{\prime}\neq 0$, we can observe that the LDOS is still localized around the rim sites for energies corresponding to the $\mathcal{E}_{\text{c}}$ band — Figure 2(b) and  2(d). The survival of imperfect CLSs is more evident from the plots of the LDOS in space Figures 4(g), 4(h), and 4(i), these density plots for $\rho(\bm{r},E)$ are obtained for an energy corresponding crossing between the upper and the central band at the K point. Here, it is possible to recognize that the LDOS still has a shape resembling the CLSs around the hub sites, with values that are much larger on the rim sites compared to the hub sites [8].
@@ II.3 Low-energy theory
In this section, we will present a low-energy, long-wavelength approximation for the Hamiltonian (2) around the K/K points. In the limit of $t^{\prime}=0$, the approximated Hamiltonian reads
@@
$\Sigma_{x}$ $=\frac{1}{\sqrt{2}}(\lambda_{1}+\lambda_{6})=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ (15a)
$\Sigma_{y}$ $=\frac{1}{\sqrt{2}}(\lambda_{2}+\lambda_{7})=\frac{1}{\sqrt{2}}\begin{pmatrix}0&-\text{i}&0\\ \text{i}&0&-\text{i}\\ 0&\text{i}&0\end{pmatrix}\!\!,$ (15b)
$\Sigma_{z}$ $=(\lambda_{3}+\sqrt{3}\lambda_{8})=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix},$ (15c)
@@
that is unitary $\Gamma^{\dagger}\Gamma=\Gamma^{2}=\mathbb{I}_{3}$ where $\mathbb{I}_{3}$ is the identity matrix of dimension 3, and anticommutes with the Hamiltonian (14): $\{\Gamma,\mathcal{H}_{0}\}=0$. This chiral symmetry operator governs the energy symmetry of the system, see Fig. 1(c)  [^1] Here the chiral symmetry is simply anticommuting with the system Hamiltonian since the flat band is at zero energy. In other lattice systems with three sites in the unit cell, such as the kagome lattice, the flat band is at finite energy. Thus they require the introduction of a generalized chiral symmetry [58, 38, 59]..
@@
We can complement the matrices (15) with two additional Gell-Mann matrices, $\lambda_{4}$ and $\lambda_{5}$, so to describe the corresponding low-energy next-nearest-neighboring hopping term $t^{\prime}$. This Hamiltonian term reads:
@@ II.4 Magnetic field properties: tight-binding approach
$\mathcal{M}$ $=\frac{t}{4}\begin{pmatrix}0&\text{e}^{-\frac{1}{2}\,\text{i}\,\pi f(6\,n_{1}-1)}&0\\ \text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}-1)}&-\frac{\Delta}{t}&\text{e}^{-\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}\\ 0&\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}&0\end{pmatrix},$ (21a)
$\mathcal{P}_{1}$ $=\begin{pmatrix}0&0&0\\ t&0&0\\ t^{\prime}\,\text{e}^{\frac{3}{2}\,\text{i}\,\pi\,f\,(2\,n_{1}+1)}&t&0\end{pmatrix},$ (21b)
$\mathcal{P}_{2}$ $=\begin{pmatrix}0&t\,\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}-1)}&t^{\prime}\\ 0&0&t\,\text{e}^{\frac{1}{2}\,\text{i}\,\pi\,f\,(6\,n_{1}+1)}\\ 0&0&0\end{pmatrix},$ (21c)
$\mathcal{P}_{3}$ $=\begin{pmatrix}0&0&0\\ 0&0&0\\ t^{\prime}\,\text{e}^{-\frac{3}{2}\,\text{i}\,\pi\,f\,(2\,n_{1}+1)}&0&0\end{pmatrix}$ (21d)
contain the magnetic hopping terms. We have omitted, for simplicity in Eq. (21a), a constant shift $\epsilon$ corresponding to the effective shift of the tight-binding model fitted over the three experimental proposals.
Typically, for the dice lattice, the magnetic flux is measured through an elementary rhombus (one-third of the unit cell surface) and normalized with respect to the $\phi_{0}=h/e$ magnetic flux quantum, that is $f=B\,\sqrt{3}\,a^{2}/(2\,\phi_{0})$. The calculation of the spectrum as a function of $f$ gives rise to the Hofstadter’s butterflies [52] presented in Figures 6 and 7. In particular, Figure 6(a) represents the ideal nearest neighbor $t=1$, $t^{\prime}=0$ case [31], whereas in Figure 6(b), we present the case of the triangular lattice corresponding to the $t=1$, $t^{\prime}=-t$ case [53]. Interestingly, for the case of the dice lattice at $f=1/2$, all the states collapse in only three energy eigenvalues. This is a consequence of the Aharonov-Bohm caging mechanism [31], for which the number of sites visited by an initially localized wave packet is limited due to destructive interferences, resulting in an extreme localization phenomenon.